### v1

# Always-Sparse Training by Growing Connections with Guided Stochastic Exploration

Abstract
The excessive computational requirements of modern artificial neural networks (ANNs) are posing limitations on the machines that can run them. Sparsification of ANNs is often motivated by time, memory and energy savings only during model inference, yielding no benefits during training. A growing body of work is now focusing on providing the benefits of model sparsification also during training. While these methods greatly improve the training efficiency, the training algorithms yielding the most accurate models still materialize the dense weights, or compute dense gradients during training. We propose an efficient, always-sparse training algorithm with excellent scaling to larger and sparser models, supported by its linear time complexity with respect to the model width during training and inference. Moreover, our guided stochastic exploration algorithm improves over the accuracy of previous sparse training methods. We evaluate our method on CIFAR-10/100 and ImageNet using ResNet, VGG, and ViT models, and compare it against a range of sparsification methods111The source code is available at: https://github.com/mikeheddes/guided-stochastic-exploration.
1 Introduction
Artificial neural networks (ANNs) are currently the most prominent machine learning method because of their unparalleled performance on a broad range of applications, including computer vision (Voulodimos et al., 2018; O’Mahony et al., 2019), natural language processing (Young et al., 2018; Otter et al., 2020), and reinforcement learning (Arulkumaran et al., 2017; Schrittwieser et al., 2020), among many others (Liu et al., 2017; Wang et al., 2020; Zhou et al., 2020). To improve their representational powers, ANNs keep increasing in size (Neyshabur et al., 2019; Kaplan et al., 2020). Recent large language models, for example, have widths surpassing units and total parameter counts over 100 billion (Brown et al., 2020; Rae et al., 2021; Chowdhery et al., 2022). However, with their increasing sizes, the memory and computational requirements to train and make inferences with these models becomes a limiting factor (Hwang, 2018; Ahmed and Wahed, 2020).
A large body of work has addressed problems arising from the immense model sizes (Reed, 1993; Gale et al., 2019; Blalock et al., 2020; Hoefler et al., 2021). Many studies look into sparsifying the model weights based on the observation that the weight distribution of a dense model is often concentrated around zero at the end of training, indicating that the majority of weights contribute little to the function being computed (Han et al., 2015). By utilizing sparse matrix representations, the model size and the number of floating-point operations (FLOPs) can be reduced dramatically. Moreover, previous work has found that for a fixed memory or parameter budget, larger sparse models outperform smaller dense models (Zhu and Gupta, 2017; Kalchbrenner et al., 2018).
Biological neural networks have also evolved to utilize sparsity, in the human brain only one in about 10 million possible connections is realized (Gerum et al., 2020). The sparsity of the brain is seen as an important property for learning efficiency (Watts and Strogatz, 1998; Bullmore and Sporns, 2009; Pessoa, 2014). To this end, sparse ANNs have shown better generalization properties (Frankle and Carbin, 2018; Chen et al., 2021), suggesting that sparsity is indeed an effective regularizer.
Early works in ANN sparsity removed connections, a process called pruning, of a trained dense model based on the magnitude of the weights (Janowsky, 1989; Ström, 1997), resulting in a more efficient model for inference. Later works improved upon this technique (Guo et al., 2016; Dong et al., 2017; Yu et al., 2018), but they all require at least the cost of training a dense model, thus yielding no efficiency benefits during training. Consequently, the resulting sparse models are limited in size by the largest dense model that can be trained on a given machine.
In light of the aforementioned limitation, the lottery ticket hypothesis (LTH), surprisingly, hypothesized that there exists a subnetwork within a dense over-parameterized model that, when trained with the same initial weights, will result in a sparse model with comparable accuracy to that of the dense model (Frankle and Carbin, 2018). However, the proposed method for finding a winning ticket within a dense model is very computationally intensive, as it requires training the dense model (typically multiple times) to obtain the subnetwork. Moreover, later work weakened the hypothesis for larger ANNs (Frankle et al., 2019). Despite this, it was still an important catalyst for new methods that aim to find the winning ticket more efficiently.
Efficient methods for finding a winning ticket can be categorized as: pruning before training and dynamic sparse training (DST). The before-training methods prune connections from a randomly initialized dense model (Lee et al., 2018; Tanaka et al., 2020). This means that they still require a machine capable of representing and computing with the dense model. In contrast, the DST methods start with a randomly initialized sparse model and change the connections dynamically throughout training while maintaining the overall sparsity (Mocanu et al., 2018; Mostafa and Wang, 2019; Evci et al., 2020). In practice, DST methods achieve higher accuracy than the pruning before training methods (see Section 4.3). The first DST method, sparse evolutionary training (SET), periodically prunes the connections with the lowest weight magnitude and grows new connections uniformly at random (Mocanu et al., 2018). While subsequent works, such as RigL and Top-KAST, have improved upon the accuracy of SET, they do so by periodically computing dense gradients and materializing the dense weights, respectively (Evci et al., 2020; Jayakumar et al., 2020).
Instead, we present a DST algorithm that is always sparse and thus scales with the number of non-zero weights, like SET, but obtains substantially higher accuracy, on par with, and at times even exceeding, that of Top-KAST and RigL. We improve upon the accuracy of previous methods using a simple, yet highly effective, method that we call guided stochastic exploration (GSE). In short, when growing new connections, our method first efficiently samples a subset of the inactive connections for exploration. It is then guided by growing the connections with the largest gradient magnitude from the sampled subset. The key insight is that when the size of the subset is on the same order as the number of active connections, that is, the weight sparsity and gradient sparsity are similar, then the accuracy of our method surpasses that of previous sparse training methods. Our method has a time complexity of with respect to the model width , where are the number of non-zero weights. When using the common Erdős–Rényi model initialization, the time complexity simplifies to just . We evaluate the accuracy of our method experimentally on CIFAR-10/100 and ImageNet using ResNet, VGG, and ViT models and compare it against a range of sparsification methods at 90%, 95%, and 98% sparsity, indicating the percentage of zero weights. Our method promises to enable the training of much larger and sparser models.
2 Related work
A wide variety of methods have been proposed that aim to reduce the size of ANNs, such as dimensionality reduction of the model weights (Jaderberg et al., 2014; Novikov et al., 2015), and weight quantization (Gupta et al., 2015; Mishra et al., 2018). However, this section only covers model sparsification methods as they are the most related to our work. Following Wang et al. (2019), the sparsification methods are categorized as: pruning after training, pruning during training, pruning before training, and dynamic sparse training.
After training
The first pruning algorithms operated on dense trained models, they pruned the connections with the smallest weight magnitude (Janowsky, 1989; Thimm and Fiesler, 1995; Ström, 1997; Han et al., 2015). This method was later generalized to first-order (Mozer and Smolensky, 1988; Karnin, 1990; Molchanov et al., 2019b, a) and second-order (LeCun et al., 1989; Hassibi and Stork, 1992; Frantar et al., 2021) Taylor polynomials of the loss with respect to the weights. These methods can be interpreted as calculating an importance score for each connection based on how its removal will effect the loss (Guo et al., 2016; Dong et al., 2017; Yu et al., 2018).
During training
Gradual pruning increases the sparsity of a model during training till the desired sparsity is reached (Zhu and Gupta, 2017; Liu et al., 2021a). Kingma et al. (2015) introduced variational dropout which adapts the dropout rate of each unit during training, Molchanov et al. (2017) showed that pruning the units with the highest dropout rate is an effective way to sparsify a model. Louizos et al. (2018) propose a method based on the reparameterization trick that allows to directly optimize the norm, which penalizes the number of non-zero weights. Alternatively, DeepHoyer is a differentiable regularizer with the same minima as the norm (Yang et al., 2019). Lastly, Powerpropagation is a sparsity promoting regularizer which raises the weights to a constant power (Schwarz et al., 2021).
Before training
The lottery ticket hypothesis (LTH) (Frankle and Carbin, 2018; Frankle et al., 2019) started a line of work that aims to find a sparse model by pruning a dense model before training (Liu et al., 2018). SNIP uses the sensitivity of each connection to the loss as the importance score of a connection (Lee et al., 2018). GraSP optimizes gradient flow to accelerate training (Wang et al., 2019), however, Lubana and Dick (2020) argue that it is better to preserve gradient flow instead. Tanaka et al. (2020) highlight a problem in the aforementioned methods: they suffer from layer collapse in high sparsity regimes, that is, during the pruning phase all the connections of a layer can be removed, making the model untrainable. They propose SynFlow, which prevents layer collapse by calculating how each connection contributes to the flow of information using a path regularizer (Neyshabur et al., 2015), similar to Lee et al. (2019).
Dynamic sparse training
The methods in this last category, including ours, start with a randomly initialized sparse model and change the connections dynamically throughout training while maintaining the same sparsity. This involves periodically pruning a fraction of active connections and growing the same number of inactive connections. SET was the first method and used magnitude pruning and random growing of connections (Mocanu et al., 2018). DeepR assigns a fixed sign to each connection at initialization and prunes those connections whose sign would change during training (Bellec et al., 2018). DSR prunes connections with the lowest global weight magnitude and uses random growing, allowing the connections to be redistributed over the layers and used in layers where they contribute the most (Mostafa and Wang, 2019). SNFS improves upon the previous methods by using an informed growing criteria: it grows the connections with the largest gradient momentum magnitude (Dettmers and Zettlemoyer, 2019). RigL makes this more efficient by only periodically calculating the dense gradients (Evci et al., 2020). However, calculating the dense gradients is still a limitation for training very large and sparse models. Top-KAST addresses this limitation by sparsely updating the dense model using only the connections with the largest weight magnitude (Jayakumar et al., 2020). While Top-KAST uses sparse gradients, it requires maintaining and updating the dense model from which the sparse model is periodically reconstructed. Moreover, we will show in Section 4.3 that Top-KAST achieves lower accuracy at higher sparsities.
3 Guided stochastic exploration
Our main contribution is an efficient dynamic sparse training algorithm which is always sparse, that is, at no point is the dense model materialized, all forward passes use sparse weights, and it exclusively calculates sparse gradients. The sparsity is maintained even when changing the connections because our method only computes gradients for a subset of the inactive connections. This subset is randomly sampled during each round of growing and pruning. The connections within the subset with the largest gradient magnitudes are then selected to grow. We call our method guided stochastic exploration (GSE), referring to the stochastic sampling process of the subset and the selection of the connections by largest gradient magnitude, which provides guidance during the exploration. Although we are not the first to present an always-sparse training method, for example, SET also achieves this, among these methods, ours obtains the highest accuracy—equal or higher than less efficient methods like RigL and Top-KAST.
We will use the following notation and definitions. With we denote the set of all possible connections, which is never materialized; is the set of non-zero or active connections whose weights are optimized with gradient descent; is a subset of inactive connections; is the set of grown connections; and is the set of pruned connections. Figure 1 illustrates the relations between the various connection sets. The solid gray area becomes the active set in the next step. The set is dotted to indicate that it is never materialized. Note that we abuse the notation of the connection sets to be per layer or global interchangeably.
GSE is designed to incorporate the strengths of both SET and RigL. While RigL typically outperforms SET, Liu et al. (2021b) showed that SET can outperform RigL when training is significantly extended. This is because RigL only explores a fraction of the connections compared to SET. The two baseline methods can be regarded as opposing extremes in their exploration strategy. On one end, SET conducts unbiased exploration, enabling it to be highly efficient. On the opposite end, RigL employs a greedy strategy by selecting connections which appear to get the largest weight magnitudes. The exploration dynamics of GSE are, in turn, determined by the interaction of the subset and the grow set. When the subset is much larger than the grow set, in the limit including all possible connections, then the exploration is nearly greedy since the sampling process has little influence on which connections are selected. On the other hand, when the sets are equal in size, then the grown connections are fully determined by the subset sampling process. The exploration strategy of GSE is thus a hybrid of random exploration guided by the selection of the largest gradient magnitude.
In the remainder of this section, we will discuss which probability distributions were considered for sampling the subset and how to efficiently sample from them in the case of fully-connected layers. In Appendix E, we discuss how our method can be applied to convolutional layers simply by interpreting them as batched fully-connected layers. We conclude this section with our complete dynamic sparse training algorithm and the time complexity.
3.1 Efficient subset sampling
For our method to be always-sparse, it is critical that we can sample a subset of the inactive connections during the grow step without explicitly representing a probability for each possible connection. Note that a connection is fully specified by the input and output units it connects. Thus, to ensure an efficient sampling process, we can sample connections at random by independently sampling from distributions over the input and output units of a layer, which induces an implicit joint probability distribution over all the connections in that layer. With this process the probability of sampling a connection is never explicitly represented, just the probability of sampling each unit is represented which requires only linear, instead of quadratic, memory and time to draw samples using the alias method (Vose, 1991).
Formally, connections are sampled from the discrete probability distributions over the input and output units of a layer : and . Since this process could sample duplicate or active connections, the set difference with the active set is computed in time using hash tables, resulting in a subset of connections whose size is bounded from above by the number of samples drawn, containing only inactive connections.
To sample the subset , perhaps the simplest distribution is uniform over the set of all possible connections . This reflects an unbiased exploration of the connections. However, the distributions f and g do not have to be uniform, it might instead be useful to bias the subset towards connections that will contribute more at decreasing the loss, that is, those that have a large gradient magnitude. To this end, we investigate two other distributions. The first distribution is called GraBo, its probabilities are proportional to an upper bound of the gradient magnitude, given as follows:
where is an all ones vector, and are the activations and gradients of the loss at the output units of the -th layer, respectively. Here, the batch size is denoted as and are the number of units in the -th layer. The second distribution is called GraEst, its probabilities are proportional to an unbiased estimate of the gradient magnitude, given as follows:
where is a vector of random signs. To compute the probabilities, both distributions can reuse the computation of the activations and the gradients during the forward and backward pass, respectively. A more detailed discussion of the distributions is provided in Appendix A. In Section 4.2, we assess which of these probability distributions is the most appropriate for sparse training.
3.2 Dynamic sparse training
Our complete dynamic sparse training algorithm starts by initializing each layer as a random bipartite graph using the Erdős–Rényi random graph generation algorithm (Erdős and Rényi, 1959). The number of active connections of a layer is given by where is a parameter that controls the sparsity of the model (Mocanu et al., 2018). This initialization ensures that the number of active connections in a layer is proportional to its width, that is, . Moreover, Evci et al. (2020) showed that the Erdős–Rényi initialization also achieves better accuracy than uniform sparsity assignment. We verify their findings in Appendix F. The weights are initialized using the procedure proposed by Evci et al. (2019) which adapts standard weight initialization to take into account the actual connectivity of each unit and scales the weight distribution for each unit accordingly.
Once the random sparse model is initialized, the weights of the active connections are optimized using stochastic gradient descent (SGD). Every training steps new connections are grown and the same number of connections are pruned. We provide pseudocode for our pruning and growing procedure in Algorithm 1, which operates on a single layer for simplicity. The maximum fraction of pruned active connections is cosine annealed from at the start of training to zero at . This ensures that there is more exploration in the early stages of training.
To determine the set of connections to grow , we start by evaluating one of the probability distributions described in Section 3.1. Then, connections are sampled from the input and output distributions, where is a hyperparameter that adjusts the number of samples. The subset contains all randomly sampled connections that are not in the active set . Now that the subset of inactive connections is determined, their gradient magnitude is computed. The grow set contains the largest gradient magnitude connections from the subset, where is the minimum between and to ensure that the grow set is not larger than the subset. The largest elements can be found in time using the introspective selection algorithm (Musser, 1997).
The pruned set , in turn, contains the active connections with the smallest weight magnitude, also determined in time. Although magnitude pruning has been observed to cause layer collapse among pruning before training methods, Tanaka et al. (2020) investigated layer collapse and found that the training process is implicitly regularizing the weight magnitudes which minimizes the risk of layer collapse for DST methods like ours. Lastly, the active set and the weights are updated to reflect the changed connections. Pseudocode for the entire training process is provided in Appendix G. It is important to stress that the dense model is not materialized at any point, only the active connections are represented, and all the operations are always sparse.
4 Experiments
At this point, in order to verify our claims, we need to determine what the minimum number of subset samples is that maintains the maximum accuracy. We answer this question in Section 4.2 and then compare the accuracy of our method against other sparsification methods in Section 4.3. We present accuracy results for ImageNet in Section 4.4. In Section 4.5, we explore the relation between model scale and accuracy for a fixed number of active connections by increasing the sparsity of larger models. Lastly, in Section 4.6, we compare the FLOPs of GSE with those of RigL.
4.1 Experiment setup
We evaluate our method and compare it to baselines on the CIFAR-10 and CIFAR-100 datasets (Krizhevsky et al., 2009), in addition to the ILSVRC-2012 ImageNet dataset (Deng et al., 2009). To establish how each method compares across model architectures we experiment with two convolutional neural networks (CNNs), ResNet-56 (He et al., 2016) and a 4 times downscaled version of VGG-16 (Simonyan and Zisserman, 2014), in addition to ViT (Dosovitskiy et al., 2020) modified for small datasets (Beyer et al., 2022) such that all models have roughly the same number of parameters. The bias and normalization parameters are kept dense since they only contribute marginally to the size and cost of a model. Similar to Wang et al. (2019), we evaluate all methods at 90%, 95%, and 98% sparsity. The high levels of sparsity emphasize the differences between the methods. In addition, in Appendix L, we show that sparse matrix formats for unstructured sparsity become more efficient for matrix multiplication from 90% sparsity compared to dense matrix formats, further motivating the need for such high levels of sparsity. We repeat each experiment three times with random initializations and report the mean and plot the 95th percentile. We adopt the implementation of the baselines based on their available code.
All experiments use the same optimization settings in order to isolate the differences in sparsification. Similar to Evci et al. (2020) and Lubana and Dick (2020), we use SGD with a momentum coefficient of 0.9, an regularization coefficient of 0.0001, and an initial learning rate of 0.1 which is dropped by a factor of 10 at epochs 80 and 120. The ViT experiments have one exception, they use an initial learning rate of 0.01. We use a batch size of 128 and train for 200 epochs. We also apply standard data augmentation to the training data, described in Appendix D. Running all the reported experiments took on the order of 2000 GPU-hours with a single NVIDIA Tesla V100 16 GB GPU per training run on an internal cluster.
The LTH procedure proposed by Frankle and Carbin (2018), denoted as Lottery in the results, uses iterative pruning with iterations at 0%, 50%, 75%, 90%, 95% and 98% sparsity. The gradual pruning method by Zhu and Gupta (2017), denoted as Gradual in the results, reaches the target sparsity at the second learning rate drop and prunes connections every 1000 steps. Following Tanaka et al. (2020), the pruning before training methods use a number of training samples equal to ten times the number of classes to prune the dense model to the target sparsity. Except for SynFlow, which uses 100 iterations with an exponential decaying sparsity schedule to prune the dense model. All the DST methods use the same update schedule: the connections are updated every 1000 steps, and similar to Evci et al. (2020), the fraction of pruned active connections is cosine annealed from 0.2 to 0.0 at the second learning rate drop. This is because Liu et al. (2021a) showed that DST methods struggle to recover from pruning when the learning rate is low. The pruning and growing procedures are applied globally for all methods, maintaining the overall sparsity. This approach generally outperforms local, layer-wise sparsity (Mostafa and Wang, 2019).
4.2 Effect of subset sample size
First, we want to determine how the size of the subset affects the accuracy. In addition, we are interested in comparing the accuracy obtained by the three distributions: uniform, gradient upper bound (GraBo), and gradient estimate (GraEst). The number of subset samples (bounding the subset size) is set proportional to the number of active connections , with and . The results are shown in the first three columns of Figure 2 and include the baseline accuracy of RigL for comparison.
We see that the accuracy of RigL is consistently matched, or even surpassed, when the number of subset samples is equal to the number of active connections, that is, . This indicates that GSE is more effective at exploring the network connections than both SET and RigL. We observe the same overall trend across all distributions, showing that the success of our method is not reliant on the specific characteristics of the sampling distribution. Moreover, the accuracy quickly peaks as , validating that it is sufficient for the subset size to be on the same order as the active set size.
We expect that as the subset size of GSE increases, the accuracy eventually drops back to the level of RigL because the largest gradient magnitude selection then becomes the dominant criteria. To this end, we experimented with increasing the number of samples to in the last column of Figure 2, note the logarithmic x-axis. The observations align precisely with our expectations, which further underscores that greedy exploration leads to inferior solutions when compared to the hybrid exploration strategy of GSE.
Interestingly, uniform sampling demonstrates strong performance while we hypothesized that the biased distributions would be beneficial for training. This shows that exploring connections which, at first, do not seem beneficial for training, can in fact be quite helpful. It would be interesting for future research to investigate exactly which training dynamics are enabled by uniform sampling. Since uniform sampling is also the most efficient distribution, we use it as our method in the following experiments.
4.3 Comparison with related work
We compare our method to a broad range of sparsification methods in Table 1, the best accuracy among the sparse training methods is bolded. The Static method denotes training a static random sparse model. The other baselines are Lottery (Frankle and Carbin, 2018), Gradual (Zhu and Gupta, 2017), SNIP (Lee et al., 2018), GraSP (Wang et al., 2019), SynFlow (Tanaka et al., 2020), SET (Mocanu et al., 2018), RigL (Evci et al., 2020), and Top-KAST (Jayakumar et al., 2020). Our results use GSE with uniform sampling and .
| Dataset | CIFAR-10 | CIFAR-100 | ||||||||||||||||
| Model | ResNet-56 | ViT | VGG-16 | ResNet-56 | ViT | VGG-16 | ||||||||||||
| Dense | 92.7 | 85.0 | 89.0 | 70.1 | 58.8 | 62.3 | ||||||||||||
| Sparsity | 90% | 95% | 98% | 90% | 95% | 98% | 90% | 95% | 98% | 90% | 95% | 98% | 90% | 95% | 98% | 90% | 95% | 98% |
| Lottery | 89.8 | 87.8 | 84.6 | 85.4 | 85.1 | 72.3 | 88.2 | 86.7 | 83.3 | 63.6 | 58.6 | 50.4 | 57.2 | 57.0 | 42.6 | 55.9 | 49.7 | 37.1 |
| Gradual | 91.3 | 89.9 | 88.0 | 83.5 | 79.3 | 72.2 | 89.1 | 88.2 | 85.8 | 67.2 | 64.4 | 57.1 | 55.6 | 53.3 | 43.6 | 60.1 | 57.0 | 44.5 |
| Static | 90.1 | 88.3 | 84.4 | 79.1 | 75.3 | 68.7 | 87.1 | 84.5 | 79.0 | 63.5 | 55.4 | 36.6 | 53.9 | 48.0 | 38.0 | 54.1 | 44.2 | 27.8 |
| SNIP | 89.6 | 87.8 | 82.9 | 75.1 | 70.4 | 63.8 | 88.3 | 86.8 | 10.0 | 62.6 | 55.5 | 41.7 | 49.7 | 44.1 | 34.6 | 52.3 | 39.1 | 1.0 |
| GraSP | 89.7 | 88.7 | 84.6 | 62.2 | 65.3 | 62.3 | 86.4 | 85.3 | 82.0 | 61.8 | 56.9 | 42.3 | 40.5 | 39.6 | 32.6 | 53.4 | 45.7 | 30.1 |
| SynFlow | 90.3 | 88.0 | 83.8 | 78.6 | 75.7 | 69.7 | 87.5 | 85.4 | 78.7 | 60.5 | 50.6 | 31.4 | 52.9 | 48.5 | 37.9 | 52.1 | 41.8 | 24.8 |
| RigL | 90.6 | 89.5 | 86.7 | 79.9 | 76.2 | 70.6 | 88.5 | 87.1 | 83.7 | 65.7 | 62.3 | 52.8 | 54.6 | 49.0 | 40.4 | 57.0 | 51.8 | 39.9 |
| Top-KAST | 89.8 | 88.1 | 85.2 | 75.0 | 68.3 | 55.0 | 86.9 | 84.8 | 80.5 | 62.5 | 58.3 | 40.4 | 51.1 | 43.3 | 28.2 | 54.4 | 47.0 | 32.2 |
| SET | 90.2 | 88.8 | 85.5 | 79.4 | 75.2 | 68.8 | 86.8 | 84.8 | 79.9 | 65.0 | 60.5 | 49.7 | 54.3 | 48.2 | 38.7 | 54.7 | 47.9 | 32.4 |
| GSE (ours) | 91.0 | 89.9 | 87.0 | 80.0 | 76.4 | 70.7 | 88.6 | 88.1 | 85.1 | 66.0 | 62.6 | 54.3 | 54.4 | 48.8 | 40.7 | 57.9 | 52.3 | 40.8 |
Among the sparse training methods (from Static downwards in Table 1) our method outperforms the other methods consistently over datasets, sparsities, and model architectures. While at 90% sparsity all the methods achieve comparable accuracy, at the extreme sparsity rate of 98%, differences between methods become more evident. At 98% sparsity our method outperforms all the before training sparsification methods by a significant margin, 4.7% on average. GSE also consistently matches or surpasses both RigL and Top-KAST, while being more efficient. For example, with VGG-16 on CIFAR-10, GSE achieves 1.4% and 4.6% higher accuracy than RigL and Top-KAST, respectively.
Gradually pruning connections during training (Gradual) can still improve the accuracy, especially in the extreme sparsity regimes. However, this comes at the cost of training the dense model, thus requiring significantly more FLOPs than our method. This observation serves as motivation for further research into dynamic sparse training methods that improve on our accuracy while preserving our efficiency advantage. Notably, our method outperforms Lottery on the CNN models, while Lottery requires multiple rounds of training.
In Appendix H, we report the test accuracy of each method at every epoch throughout training, showing that our method trains as fast as the other sparse training methods. Additionally, in Appendix M, we provide plots of the layer-wise sparsity at the end of training for each combination of method, model, and sparsity. This information provides insight into how each method distributes the available connections over the layers of a model.
4.4 ImageNet
| Method | Accuracy |
| Dense | 76.8±0.09 |
| Lottery | 69.4±0.03 |
| Gradual | 73.9 |
| Static | 67.7±0.12 |
| SNIP | 67.2±0.12 |
| GraSP | 68.1 |
| SET | 69.6±0.23 |
| DSR | 71.6 |
| SNFS | 72.9±0.06 |
| RigL | 73.0±0.04 |
| Top-KAST | 73.0 |
| GSE (ours) | 73.2±0.07 |
To determine the efficacy and robustness of GSE on large models and datasets, in this third experiment, we compare the accuracy of ResNet-50 trained on ImageNet at 90% sparsity. In Table 2, the accuracy obtained with GSE is compared against the baselines from the preceding section, with the additions of DSR (Mostafa and Wang, 2019) and SNFS (Dettmers and Zettlemoyer, 2019). Since this is a common benchmark among sparse training methods, the accuracy results of the baselines were obtained from prior publications.
Our experiments use the aforementioned hyperparameters with the following exceptions: we train for 100 epochs and use learning rate warm-up over the initial 5 epochs, gradually reaching a value of 0.1; subsequently, the learning rate is reduced by a factor of 10 every 30 epochs; and we use label smoothing with a coefficient of 0.1. Our results use GSE with uniform sampling and . The results in Table 2 align with those for CIFAR, and further underscore GSE’s superior performance over other sparse training methods. Note also that GSE again outperforms less efficient methods such as SNFS, RigL, and Top-KAST. These results collectively reinforce that GSE is robust and able to scale to large models and datasets.
4.5 Model scaling
In this experiment, we investigate how the size of a model affects the accuracy when the number of active connections is kept constant. We scale the number of filters in convolution layers and units in fully-connected layers to obtain a wider model. The results, presented in Figure 3, show that training wider CNN models translates to significantly increased accuracy, 6.5% on average at 4 times wider. These findings support those from Zhu and Gupta (2017) and Kalchbrenner et al. (2018), however, our method enables a sparse model to be trained whose dense version would be too computationally intensive.
The same trend does not hold for ViT, as its accuracy decreases with increasing scale. Appendix J reveals that the accuracy increases slower for larger ViT models, while for CNNs the accuracy increases faster as they get larger. Moreover, the accuracy of larger dense ViT models plateaus at 62.2%, as shown in Appendix K, indicating that this phenomenon is independent of our method. ViT models are known to require large amounts of training data to achieve optimal accuracy (Dosovitskiy et al., 2020), we conjecture that this contributes to the trend observed with ViT. In Appendix I, we observe that both ViT and CNN models do benefit considerably from extended training, with a 5.3% average increase in accuracy for training twice as long.
4.6 Floating-point operations
Lastly, to determine the efficiency of GSE, we compare its FLOPs to those of RigL in Figure 4. The FLOPs are obtained for ResNet-50 on ImageNet using the method described by Evci et al. (2020), with . The results show that at 99% sparsity GSE uses 11.8% fewer FLOPs than RigL, at higher sparsities GSE uses considerably fewer FLOPs. For illustration, the right plot extends the sparsity to that of various animal brains which underscores the significance of our always sparse training method, as RigL uses times more FLOPs to train a model with the same sparsity as the cat brain.
4.7 Limitations
While we experiment on a range of sparsities, datasets, models, and baseline methods, we acknowledge that further exploration would enhance the robustness of our findings. Following previous work, our experiments focus on the domain of computer vision, and although we demonstrate the effectiveness of our approach within this domain, it remains to be seen how our results generalize to other domains such as natural language processing and reinforcement learning, among others. Despite this, we believe that our proposed always-sparse training algorithm, supported by its linear time complexity with respect to the model width during training and inference, represents a significant advancement. Furthermore, our method outperforms existing sparse training algorithms in terms of accuracy, as demonstrated on the CIFAR and ImageNet datasets using ResNet, VGG, and ViT models, alongside a comprehensive comparison with a range of sparsification techniques.
Secondly, our accuracy results were obtained by simulating sparse networks using binary masks. This is standard practice among sparse training methods because support for sparsity in machine learning frameworks is still in the early days, making it challenging to train models using sparse matrix formats for the weights. We note that there is nothing about our method that prevents it from being accelerated on GPUs, yet the significant engineering required to develop an optimized GPU kernel falls outside the scope of this work. To ensure that our claims are still validated, we compare the FLOPs of our method with those of RigL in Section 4.6. Moreover, we present an unstructured sparse matrix multiplication benchmark in Appendix L, which shows that a 10 times reduction in execution time is achieved at 99% sparsity, compared to dense matrix multiplication. Thus, evidently showing that real speedups can be achieved, even with unstructured sparsity.
5 Conclusion
We presented an always-sparse training algorithm using guided stochastic exploration of the connections which has a linear time complexity with respect to the model width and improves upon the accuracy of previous sparse training methods. This is achieved by only evaluating gradients for a subset of randomly sampled connections when changing the connections. The three distributions for sampling the subset—uniform, an upper bound of the gradient magnitude, and an estimate of the gradient magnitude—all showed similar trends. However, uniform consistently achieves among the highest accuracy and is also the most efficient. We therefore conclude that the uniform distribution is the most appropriate for sparse training. Interestingly, we showed that training larger and sparser CNN models results in higher accuracy for the same number of active connections. Lastly, the number of floating-point operations of our method decrease considerably faster than those for RigL as the sparsity increases. This observation underscores the efficiency and scalability of our method, which is promising for training increasingly larger and sparser ANNs.
Appendix A Connection sampling distributions
In this section we provide a more detailed discussion on the distributions which were considered for sampling the subset of inactive connections.
A.1 Gradient magnitude upper bound (GraBo)
The gradient magnitude for a single connection can be expressed as follows:
where and are the activations and gradients of the loss at the output units of the -th layer, respectively. When the batch size is one, the sum disappears, simplifying the gradient magnitude to a rank one matrix which can be used to sample connections efficiently, inducing a joint probability distribution that is proportional to the gradient magnitude only when the batch size is one. In practice, however, training samples come in mini batches to reduce the variance of the gradient. Therefore, we experiment with sampling the connections proportionally to the following upper bound of the gradient magnitude instead, which enables efficient sampling, even with mini batches:
The proof for Equation 2 uses the triangle inequality and is provided in Appendix B. This upper bound has the property that it becomes an equality when . Connections are then sampled from the probability distributions in Equation 3. The implementation of this distribution does not require any modifications to the model, and only needs a single forward and backward pass to evaluate.
A.2 Gradient magnitude estimate (GraEst)
The reason we use a bound on the absolute gradient is that in general the sum of products cannot be written as the product of sums, that is, . However, it is possible to satisfy this equality in expectation. This was shown by Alon et al. (1996) who used it to estimate the second frequency moment, or the norm of a vector. Their work is called the AMS sketch, AMS is an initialism derived from the last names of the authors. They then showed that this also works in the more general case to estimate inner products between vectors (Alon et al., 1999). This is achieved by assigning each index of the vector a random sign, i.e. , resulting in the following equality:
We can use their finding to construct probability distributions and in Equation 5 such that their induced joint probability distribution is proportional to the gradient magnitude in expectation. This distribution can also be computed in a single forward and backward pass and requires no modifications to the model, apart from introducing a mechanism to sample the random signs .
Appendix B Gradient magnitude upper bound proof
Lemma B.1.
The gradient magnitude of the loss with respect to the parameters of a fully-connected layer has the following upper bound:
where , , and are the weight matrix, activations and gradients of the loss at the output units of the -th layer, respectively. The number of units in the -th layer is denoted by , and is the batch size.
Proof.
The proof starts with the definition of the gradient of the fully-connected layer, which is calculated during back propagation as the activations of the previous layer multiplied with the back propagated gradient of the output units of the layer. We then write this definition for a single connection in the weight matrix and use the triangle inequality to specify the first upper bound. In Equation 8 we rewrite the upper bound of Equation 7 as all the cross products between the batch dimensions minus all the non-identical cross terms. Note that all the non-identical cross terms are positive because of the absolute, therefore we get another upper bound by removing the non-identical cross terms which we then write in matrix notation in Equation 9.
Appendix C Data normalization
We normalize all the training and test data to have zero mean and a standard deviation of one. The dataset statistics that we used are specified below, where is the mean and the standard deviation. The values correspond to the red, green, and blue color channels of the image, respectively.
CIFAR-10
CIFAR-100
ImageNet
Appendix D Data augmentation
We applied standard data augmentation as part of our training data pipeline. The specific augmentation techniques are specified below. We apply the data normalization described in Appendix C after the augmentation transformations.
CIFAR-10
We pad the image with 4 black pixels on all sides and randomly crop a square of 32 by 32 pixels. We then perform a random horizontal flip of the crop with a probability of 0.5.
CIFAR-100
We pad the image with 4 black pixels on all sides and randomly crop a square of 32 by 32 pixels. We then perform a random horizontal flip of the crop with a probability of 0.5.
ImageNet
We randomly crop a square of 224 by 224 pixels. We then perform a random horizontal flip of the crop with a probability of 0.5. For the test set, we resize the images to be 256 pixels followed by a square center crop of 224 pixels.
Appendix E Convolutional layer subset sampling
In this section we describe how our growing and pruning algorithm can be applied to convolutional layers. While a weight in a fully-connected layer is specified by the input and output units that it connects, a weight in a 2D convolutional layer is specified by the input channel, its and coordinate on the filter, and the output channel. It thus seems that it would require four discrete probability distributions to sample a weight of a 2D convolutional layer, instead of two for fully-connected layers. However, one can alternatively view a convolutional layer as applying a fully-connected layer on multiple subsections (patches) of the input, flattening the input channels and filters makes the weights of a convolutional layer identical to a fully-connected layer. The multiple subsections of the input can then simply be treated as additional batches. With this realization all the methods that we explain throughout the paper for fully-connected layers directly apply to convolutional layers.
Appendix F Sparse graph initialization comparison
In Figure 5, we plot the accuracy obtained by assigning the sparsity uniformly over all the layers compared to using the Erdős–Rényi initialization for SET, RigL, and our method. Note that Ours Uniform (Erdős–Rényi) denotes uniform sampling of the subset connections and the Erdős–Rényi initialization. The Erdős–Rényi initialization obtains better accuracy while ensuring that the number of active connections per layer is proportional to the layer width.
Appendix G Pseudocode
We provide pseudocode for our complete dynamic sparse training process in the case of the GraBo or GraEst distributions in Algorithm 2. The uniform distribution is simpler as it does not require any aggregation of activity. For simplification, the pseudocode assumes that the prune-grow step only uses a single batch of training samples, allowing the aggregation for the probability distributions f and g to be computed in the same pass as the generation of the various connection sets. This is also the scenario we tested with in all our experiments.
Appendix H Training progress
In Figure 6, we provide plots for the test accuracy during training for each baseline method and our method. It can be seen that our method consistently achieves among the highest test accuracy throughout training compared to the sparse training methods. The training progress of Gradual stands out because its accuracy increases the most at first, this is because it starts out as a dense model while it gradually increases the model sparsity, as a result, the accuracy goes down towards the second learning rate drop when the target sparsity is reached.
Appendix I Extended training
The following extended training results are for CIFAR-100 at 98% sparsity with 1x, 1.5x, and 2x the number of training epochs. The learning rate drops and the final prune-grow step are also scaled by the same amount, all other hyperparameters stay the same. It is clear that all models benefit significantly from extended training. The ViT model in particular gains 7% in accuracy by training for twice as long.
| Epochs | 200 (1x) | 300 (1.5x) | 400 (2x) |
| ResNet-56 | 54.3±0.5 | 56.4±0.3 | 57.9±0.1 |
| VGG-16 | 40.8±0.3 | 44.1±0.4 | 45.9±0.2 |
| ViT | 40.7±1.0 | 45.3±1.1 | 47.8±0.5 |
Appendix J Model scaling training progress
In Figure 7, we provide plots for the test and training accuracy during training for our method on CIFAR-100 while increasing the width of the model by 1x, 2x, 3x, and 4x. The number of active connections is kept constant between all experiments, see Section 4.5 for details. In the CNN models, ResNet-56 and VGG-16, the increase in width causes the models to train significantly faster and obtain higher accuracy. For the vision transformer (ViT) model, the trend is reversed, wider models train slower and obtain lower accuracy.
Appendix K Scaling dense ViT
To further investigate the behavior of ViT, which decreases in accuracy when using larger models with the same number of activate parameters, as discussed in Section 4.5 and Appendix J. In Table 4, we report the accuracy of ViT for increasing model width but kept dense, thus increasing the active connections count. We see the accuracy quickly plateaus at 62.2%, gaining only 2% accuracy while even the sparse CNNs gained 6.5% on average at 4 times wider. This indicates that the decreasing accuracy of ViT reported in Section 4.5 is not due to our sparse training method as even the dense model does not benefit much from increased scale.
| Model width | 1x | 2x | 3x | 4x |
| Accuracy | 60.3±0.6 | 62.5±0.5 | 62.2±0.4 | 62.2±0.1 |
Appendix L Sparse matrix multiply benchmark
In Figure 8, we show the execution time of the sparse times dense matrix-matrix multiply on two execution devices: Intel Xeon Gold 6148 CPU at 2.40GHz and NVIDIA Tesla V100 16GB. The sparse matrix is a square matrix of size units times units, representing an unstructured sparse random weight matrix. The dense matrix is of size units times 128 and represents a batch of layer inputs. We compare three storage formats, namely: dense, where all the zeros are explicitly represented; Coordinate format (COO), which represents only the nonzero elements and their coordinates; and Compressed Sparse Row (CSR), which is similar to COO but compresses the row indices. Each data point in Figure 8 is the average of 10 measurements.
The results show that the CSR format generally outperforms the COO format. Moreover, the execution time of the CSR format improves upon the dense execution time starting around 90% sparsity. At 99% sparsity, the CSR format is up to an order of magnitude faster than the dense format. These results are similar between the two execution devices, although the NVIDIA Tesla V100 16GB is up to two orders of magnitude faster than the Intel Xeon Gold 6148 CPU at 2.40GHz.
Appendix M Layer-wise sparsity
In Figure 9, we show the sparsity of each layer, in terms of compression, obtained by each sparsification methods at the end of training. In the plots the CIFAR-10 and CIFAR-100 results are averaged. These plots can be used to gain insight into how each sparsification method distributes the active connections over the models.

### v2

@@
 Moreover, our guided stochastic exploration algorithm improves over the accuracy of previous sparse training methods.
-We evaluate our method on CIFAR-10/100 and ImageNet using ResNet, VGG, and ViT models, and compare it against a range of sparsification methods111The source code is available at: https://github.com/mikeheddes/guided-stochastic-exploration.
-1 Introduction
-Artificial neural networks (ANNs) are currently the most prominent machine learning method because of their unparalleled performance on a broad range of applications, including computer vision (Voulodimos et al., 2018; O’Mahony et al., 2019), natural language processing (Young et al., 2018; Otter et al., 2020), and reinforcement learning (Arulkumaran et al., 2017; Schrittwieser et al., 2020), among many others (Liu et al., 2017; Wang et al., 2020; Zhou et al., 2020).
-To improve their representational powers, ANNs keep increasing in size (Neyshabur et al., 2019; Kaplan et al., 2020).
-Recent large language models, for example, have widths surpassing units and total parameter counts over 100 billion (Brown et al., 2020; Rae et al., 2021; Chowdhery et al., 2022).
-However, with their increasing sizes, the memory and computational requirements to train and make inferences with these models becomes a limiting factor (Hwang, 2018; Ahmed and Wahed, 2020).
-A large body of work has addressed problems arising from the immense model sizes (Reed, 1993; Gale et al., 2019; Blalock et al., 2020; Hoefler et al., 2021).
-Many studies look into sparsifying the model weights based on the observation that the weight distribution of a dense model is often concentrated around zero at the end of training, indicating that the majority of weights contribute little to the function being computed (Han et al., 2015).
-By utilizing sparse matrix representations, the model size and the number of floating-point operations (FLOPs) can be reduced dramatically.
-Moreover, previous work has found that for a fixed memory or parameter budget, larger sparse models outperform smaller dense models (Zhu and Gupta, 2017; Kalchbrenner et al., 2018).
-Biological neural networks have also evolved to utilize sparsity, in the human brain only one in about 10 million possible connections is realized (Gerum et al., 2020).
-The sparsity of the brain is seen as an important property for learning efficiency (Watts and Strogatz, 1998; Bullmore and Sporns, 2009; Pessoa, 2014).
-To this end, sparse ANNs have shown better generalization properties (Frankle and Carbin, 2018; Chen et al., 2021), suggesting that sparsity is indeed an effective regularizer.
-Early works in ANN sparsity removed connections, a process called pruning, of a trained dense model based on the magnitude of the weights (Janowsky, 1989; Ström, 1997), resulting in a more efficient model for inference.
-Later works improved upon this technique (Guo et al., 2016; Dong et al., 2017; Yu et al., 2018), but they all require at least the cost of training a dense model, thus yielding no efficiency benefits during training.
+We evaluate our method on the CIFAR-10/100 and ImageNet classification tasks using ResNet, VGG, and ViT models, and compare it against a range of sparsification methods333The source code is available at:
+https://github.com/mikeheddes/guided-stochastic-exploration.
+Index Terms:
+dynamic sparse training, neural network pruning, sparse neural networksI Introduction
+Artificial neural networks (ANNs) are currently the most prominent machine learning method due to their unparalleled performance in a wide range of applications, including computer vision [1, 2], natural language processing [3, 4], and reinforcement learning [5, 6], among many others [7, 8, 9].
+To improve their representational powers, ANNs continue to increase in size [10, 11].
+Recent large language models, for example, have widths exceeding units and total parameter counts over 100 billion [12, 13, 14].
+However, with their increasing sizes, the memory and computational requirements to train and make inferences with these models become a limiting factor [15, 16].
+A large body of work has addressed problems that arise from the immense model sizes [17, 18, 19, 20].
+Many studies look into sparsifying the model weights based on the observation that the weight distribution of a dense model is often concentrated around zero at the end of training, indicating that the majority of weights contribute little to the function being computed [21].
+Using sparse matrix representations, the model size and the number of floating-point operations (FLOPs) can be dramatically reduced.
+Moreover, previous work has found that for a fixed memory or parameter budget, larger sparse models outperform smaller dense models [22, 23].
+Early works in ANN sparsity removed connections, a process called pruning, of a trained dense model based on the magnitude of the weights [24, 25], resulting in a more efficient model for inference.
+Later works improved upon this technique [26, 27, 28], but all require at least the cost of training a dense model, thus providing no efficiency benefits during training.
 Consequently, the resulting sparse models are limited in size by the largest dense model that can be trained on a given machine.
-In light of the aforementioned limitation, the lottery ticket hypothesis (LTH), surprisingly, hypothesized that there exists a subnetwork within a dense over-parameterized model that, when trained with the same initial weights, will result in a sparse model with comparable accuracy to that of the dense model (Frankle and Carbin, 2018).
+In light of the aforementioned limitation, the lottery ticket hypothesis (LTH), surprisingly, hypothesized that there exists a subnetwork within a dense over-parameterized model that, when trained with the same initial weights, will result in a sparse model with accuracy comparable to that of the dense model [29].
 However, the proposed method for finding a winning ticket within a dense model is very computationally intensive, as it requires training the dense model (typically multiple times) to obtain the subnetwork.
-Moreover, later work weakened the hypothesis for larger ANNs (Frankle et al., 2019).
+Moreover, later work weakened the hypothesis for larger ANNs [30].
 Despite this, it was still an important catalyst for new methods that aim to find the winning ticket more efficiently.
-Efficient methods for finding a winning ticket can be categorized as: pruning before training and dynamic sparse training (DST).
-The before-training methods prune connections from a randomly initialized dense model (Lee et al., 2018; Tanaka et al., 2020).
+Efficient methods for finding a winning ticket can be categorized as pruning before training and dynamic sparse training (DST).
+The before-training methods prune connections from a randomly initialized dense model [31, 32].
 This means that they still require a machine capable of representing and computing with the dense model.
-In contrast, the DST methods start with a randomly initialized sparse model and change the connections dynamically throughout training while maintaining the overall sparsity (Mocanu et al., 2018; Mostafa and Wang, 2019; Evci et al., 2020).
-In practice, DST methods achieve higher accuracy than the pruning before training methods (see Section 4.3).
-The first DST method, sparse evolutionary training (SET), periodically prunes the connections with the lowest weight magnitude and grows new connections uniformly at random (Mocanu et al., 2018).
-While subsequent works, such as RigL and Top-KAST, have improved upon the accuracy of SET, they do so by periodically computing dense gradients and materializing the dense weights, respectively (Evci et al., 2020; Jayakumar et al., 2020).
-Instead, we present a DST algorithm that is always sparse and thus scales with the number of non-zero weights, like SET, but obtains substantially higher accuracy, on par with, and at times even exceeding, that of Top-KAST and RigL.
-We improve upon the accuracy of previous methods using a simple, yet highly effective, method that we call guided stochastic exploration (GSE).
-In short, when growing new connections, our method first efficiently samples a subset of the inactive connections for exploration.
+In contrast, DST methods start with a randomly initialized sparse model and dynamically change the connections throughout training while maintaining the overall sparsity [33, 34, 35], as illustrated in Figure 1.
+In practice, DST methods achieve higher accuracy than pruning before training methods (see Section IV-B).
+The first DST method, sparse evolutionary training (SET), periodically prunes the connections with the lowest weight magnitude and grows new connections uniformly at random [33].
+Subsequent work, such as RigL and Top-KAST [35, 36], has improved the accuracy of SET.
+However, RigL periodically computes dense gradients, and while Top-KAST always computes sparse gradients, it still materializes dense weights during training.
+Instead, we present a DST algorithm that is always sparse, i.e., both the gradients and the weights are sparse at all times.
+The training and inference cost of our method therefore scales with the number of non-zero weights, like SET, but obtains substantially higher accuracy, on par with, and at times even exceeding, that of Top-KAST and RigL.
+We improve the accuracy of previous methods using a simple, yet effective method that we call guided stochastic exploration (GSE).
+In short, when new connections are added, our method first efficiently samples a subset of inactive connections for exploration.
 It is then guided by growing the connections with the largest gradient magnitude from the sampled subset.
 The key insight is that when the size of the subset is on the same order as the number of active connections, that is, the weight sparsity and gradient sparsity are similar, then the accuracy of our method surpasses that of previous sparse training methods.
-Our method has a time complexity of with respect to the model width , where are the number of non-zero weights.
-When using the common Erdős–Rényi model initialization, the time complexity simplifies to just .
-We evaluate the accuracy of our method experimentally on CIFAR-10/100 and ImageNet using ResNet, VGG, and ViT models and compare it against a range of sparsification methods at 90%, 95%, and 98% sparsity, indicating the percentage of zero weights.
-Our method promises to enable the training of much larger and sparser models.
-2 Related work
-A wide variety of methods have been proposed that aim to reduce the size of ANNs, such as dimensionality reduction of the model weights (Jaderberg et al., 2014; Novikov et al., 2015), and weight quantization (Gupta et al., 2015; Mishra et al., 2018).
-However, this section only covers model sparsification methods as they are the most related to our work.
+Our method has a time complexity of with respect to the width of the model , where is the number of non-zero weights.
+When using the Erdős–Rényi sparse model initialization [33], the time complexity is reduced to only .
+We evaluated the classification accuracy of our method on the CIFAR-10/100 and ImageNet datasets using ResNet, VGG, and ViT models and compared it with a range of sparsification methods at 90%, 95%, and 98% sparsity, indicating the percentage of zero weights.
+Our method promises to enable the training of larger and sparser models.
+II Related work
+A wide variety of methods have been proposed that aim to reduce the size of ANNs, such as dimensionality reduction of the model weights [37, 38], and weight quantization [39, 40].
+However, this section specifically covers model sparsification methods as they are the most related to our work.
 Following Wang et al.
-(2019), the sparsification methods are categorized as: pruning after training, pruning during training, pruning before training, and dynamic sparse training.
+[41], the sparsification methods are categorized as: pruning after training, pruning during training, pruning before training, and dynamic sparse training.
 After training
-The first pruning algorithms operated on dense trained models, they pruned the connections with the smallest weight magnitude (Janowsky, 1989; Thimm and Fiesler, 1995; Ström, 1997; Han et al., 2015).
-This method was later generalized to first-order (Mozer and Smolensky, 1988; Karnin, 1990; Molchanov et al., 2019b, a) and second-order (LeCun et al., 1989; Hassibi and Stork, 1992; Frantar et al., 2021) Taylor polynomials of the loss with respect to the weights.
-These methods can be interpreted as calculating an importance score for each connection based on how its removal will effect the loss (Guo et al., 2016; Dong et al., 2017; Yu et al., 2018).
+The first pruning algorithms operated on dense trained models, they prune the connections with the smallest weight magnitude [24, 42, 25, 21].
+This method was later generalized to first-order [43, 44, 45, 46] and second-order [47, 48, 49] Taylor polynomials of the loss with respect to the weights.
+These methods can be interpreted as calculating an importance score for each connection based on how its removal will affect the loss [26, 27, 28].
 During training
-Gradual pruning increases the sparsity of a model during training till the desired sparsity is reached (Zhu and Gupta, 2017; Liu et al., 2021a).
+Gradual pruning increases the sparsity of a model during training untill the desired sparsity is reached [22, 50].
 Kingma et al.
-(2015) introduced variational dropout which adapts the dropout rate of each unit during training, Molchanov et al.
-(2017) showed that pruning the units with the highest dropout rate is an effective way to sparsify a model.
+[51] introduced variational dropout which adapts the dropout rate of each unit during training, Molchanov et al.
+[52] showed that pruning the units with the highest dropout rate is an effective way to sparsify a model.
 Louizos et al.
-(2018) propose a method based on the reparameterization trick that allows to directly optimize the norm, which penalizes the number of non-zero weights.
-Alternatively, DeepHoyer is a differentiable regularizer with the same minima as the norm (Yang et al., 2019).
-Lastly, Powerpropagation is a sparsity promoting regularizer which raises the weights to a constant power (Schwarz et al., 2021).
+[53] propose a method based on the reparameterization trick that allows to directly optimize the norm, which penalizes the number of non-zero weights.
+Alternatively, DeepHoyer is a differentiable regularizer with the same minima as the norm [54].
+Lastly, Powerpropagation is a sparsity-promoting regularizer that raises weights to a constant power [55].
 Before training
-The lottery ticket hypothesis (LTH) (Frankle and Carbin, 2018; Frankle et al., 2019) started a line of work that aims to find a sparse model by pruning a dense model before training (Liu et al., 2018).
-SNIP uses the sensitivity of each connection to the loss as the importance score of a connection (Lee et al., 2018).
-GraSP optimizes gradient flow to accelerate training (Wang et al., 2019), however, Lubana and Dick (2020) argue that it is better to preserve gradient flow instead.
+The lottery ticket hypothesis (LTH) [29, 30] started a line of work that aims to find a sparse model by pruning a dense model before training [56].
+SNIP uses the sensitivity of each connection to the loss as the importance score of a connection [31].
+GraSP optimizes gradient flow to accelerate training [41], however, Lubana et al.
+[57] argue that it is better to preserve gradient flow instead.
+Gradient flow refers to the norm of the gradients after pruning.
 Tanaka et al.
-(2020) highlight a problem in the aforementioned methods: they suffer from layer collapse in high sparsity regimes, that is, during the pruning phase all the connections of a layer can be removed, making the model untrainable.
-They propose SynFlow, which prevents layer collapse by calculating how each connection contributes to the flow of information using a path regularizer (Neyshabur et al., 2015), similar to Lee et al.
-(2019).
+[32] highlight a problem in the aforementioned methods: they suffer from layer collapse in high-sparsity regimes, that is, during the pruning phase all the connections of a layer can be removed, making the model untrainable.
+They propose SynFlow, which prevents layer collapse by calculating how each connection contributes to the flow of information using a path regularizer [58], similar to Lee et al.
+[59].
 Dynamic sparse training
-The methods in this last category, including ours, start with a randomly initialized sparse model and change the connections dynamically throughout training while maintaining the same sparsity.
+The methods in this last category, including ours, start with a randomly initialized sparse model and dynamically change the connections throughout training while maintaining the same sparsity.
 This involves periodically pruning a fraction of active connections and growing the same number of inactive connections.
-SET was the first method and used magnitude pruning and random growing of connections (Mocanu et al., 2018).
-DeepR assigns a fixed sign to each connection at initialization and prunes those connections whose sign would change during training (Bellec et al., 2018).
-DSR prunes connections with the lowest global weight magnitude and uses random growing, allowing the connections to be redistributed over the layers and used in layers where they contribute the most (Mostafa and Wang, 2019).
-SNFS improves upon the previous methods by using an informed growing criteria: it grows the connections with the largest gradient momentum magnitude (Dettmers and Zettlemoyer, 2019).
-RigL makes this more efficient by only periodically calculating the dense gradients (Evci et al., 2020).
+This process is illustrated in Figure 1.
+SET was the first method and used magnitude pruning and random growing of connections [33].
+DeepR assigns a fixed sign to each connection at initialization and prunes those connections whose sign would change during training [60].
+DSR prunes connections with the lowest global weight magnitude and uses random growing, allowing the connections to be redistributed over the layers and used in layers where they contribute the most [34].
+SNFS improves upon the previous methods by using an informed growing criteria: it grows the connections with the largest gradient momentum magnitude [61].
+RigL makes this more efficient by periodically calculating the dense gradients [35].
 However, calculating the dense gradients is still a limitation for training very large and sparse models.
-Top-KAST addresses this limitation by sparsely updating the dense model using only the connections with the largest weight magnitude (Jayakumar et al., 2020).
-While Top-KAST uses sparse gradients, it requires maintaining and updating the dense model from which the sparse model is periodically reconstructed.
-Moreover, we will show in Section 4.3 that Top-KAST achieves lower accuracy at higher sparsities.
-3 Guided stochastic exploration
+Top-KAST addresses this limitation by sparsely updating the dense model using only the connections with the largest weight magnitude [36].
+Although Top-KAST uses sparse gradients, it requires maintaining and updating the dense model from which the sparse model is periodically reconstructed.
+Moreover, we will show in Section IV-B that Top-KAST achieves lower accuracy at higher sparsities.
+III Guided stochastic exploration
 Our main contribution is an efficient dynamic sparse training algorithm which is always sparse, that is, at no point is the dense model materialized, all forward passes use sparse weights, and it exclusively calculates sparse gradients.
-The sparsity is maintained even when changing the connections because our method only computes gradients for a subset of the inactive connections.
-This subset is randomly sampled during each round of growing and pruning.
+The sparsity is maintained even while changing the connections because our method only computes gradients for a subset of the inactive connections.
+This subset is randomly sampled using an efficient sampling procedure during each round of growing and pruning.
 The connections within the subset with the largest gradient magnitudes are then selected to grow.
@@
 We will use the following notation and definitions.
-With we denote the set of all possible connections, which is never materialized; is the set of non-zero or active connections whose weights are optimized with gradient descent; is a subset of inactive connections; is the set of grown connections; and is the set of pruned connections.
+With we denote the set of all possible connections, which is never materialized; is the set of active connections whose weights are optimized with gradient descent, while the inactive connections have an implicit weight of zero; is a subset of inactive connections; is the set of grown connections; and is the set of pruned connections.
 Figure 2 illustrates the relations between the various connection sets.
-The solid gray area becomes the active set in the next step.
-The set is dotted to indicate that it is never materialized.
-Note that we abuse the notation of the connection sets to be per layer or global interchangeably.
+Note that we abuse the notation of connection sets to be per layer or global interchangeably.
 GSE is designed to incorporate the strengths of both SET and RigL.
 While RigL typically outperforms SET, Liu et al.
-(2021b) showed that SET can outperform RigL when training is significantly extended.
+[62] showed that SET can outperform RigL when training is significantly extended.
 This is because RigL only explores a fraction of the connections compared to SET.
@@
 The exploration dynamics of GSE are, in turn, determined by the interaction of the subset and the grow set.
-When the subset is much larger than the grow set, in the limit including all possible connections, then the exploration is nearly greedy since the sampling process has little influence on which connections are selected.
+When the subset is much larger than the grow set, in the limit including all possible connections, then the exploration is nearly greedy since the sampling process has little influence over which connections are selected.
 On the other hand, when the sets are equal in size, then the grown connections are fully determined by the subset sampling process.
 The exploration strategy of GSE is thus a hybrid of random exploration guided by the selection of the largest gradient magnitude.
-In the remainder of this section, we will discuss which probability distributions were considered for sampling the subset and how to efficiently sample from them in the case of fully-connected layers.
-In Appendix E, we discuss how our method can be applied to convolutional layers simply by interpreting them as batched fully-connected layers.
-We conclude this section with our complete dynamic sparse training algorithm and the time complexity.
-3.1 Efficient subset sampling
+In the remainder of this section, we will discuss which probability distributions were considered for sampling the subset and how to efficiently sample from them in the case of feedforward layers.
+Our method naturally applies to convolutional layers by interpreting them as batched feedforward layers.
+See Appendix -E for an extended discussion.
+We conclude this section with our complete dynamic sparse training algorithm and its time complexity.
+III-A Efficient subset sampling
 For our method to be always-sparse, it is critical that we can sample a subset of the inactive connections during the grow step without explicitly representing a probability for each possible connection.
@@
 Thus, to ensure an efficient sampling process, we can sample connections at random by independently sampling from distributions over the input and output units of a layer, which induces an implicit joint probability distribution over all the connections in that layer.
-With this process the probability of sampling a connection is never explicitly represented, just the probability of sampling each unit is represented which requires only linear, instead of quadratic, memory and time to draw samples using the alias method (Vose, 1991).
-Formally, connections are sampled from the discrete probability distributions over the input and output units of a layer : and .
-Since this process could sample duplicate or active connections, the set difference with the active set is computed in time using hash tables, resulting in a subset of connections whose size is bounded from above by the number of samples drawn, containing only inactive connections.
+With this process the probability of sampling a connection is never explicitly represented, just the probability of sampling each unit is represented which requires only linear, instead of quadratic, memory and time to draw samples using the alias method [63].
+Formally, connections are sampled from the discrete probability distributions over the input and output units of the -th layer: and .
+Since this process could sample duplicate or active connections, the set difference with the active set is computed in time using hash tables, resulting in a subset of connections whose size is upper bounded by the number of samples drawn, containing only inactive connections.
 To sample the subset , perhaps the simplest distribution is uniform over the set of all possible connections .
@@
 The first distribution is called GraBo, its probabilities are proportional to an upper bound of the gradient magnitude, given as follows:
-where is an all ones vector, and are the activations and gradients of the loss at the output units of the -th layer, respectively.
-Here, the batch size is denoted as and are the number of units in the -th layer.
+where is an all-ones vector, and are the activations and gradients of the loss at the output units of the -th layer, respectively.
+Here, the batch size is denoted as and is the number of units in the -th layer.
 The second distribution is called GraEst, its probabilities are proportional to an unbiased estimate of the gradient magnitude, given as follows:
@@
 To compute the probabilities, both distributions can reuse the computation of the activations and the gradients during the forward and backward pass, respectively.
-A more detailed discussion of the distributions is provided in Appendix A.
-In Section 4.2, we assess which of these probability distributions is the most appropriate for sparse training.
-3.2 Dynamic sparse training
-Our complete dynamic sparse training algorithm starts by initializing each layer as a random bipartite graph using the Erdős–Rényi random graph generation algorithm (Erdős and Rényi, 1959).
-The number of active connections of a layer is given by where is a parameter that controls the sparsity of the model (Mocanu et al., 2018).
+A more detailed discussion of the distributions is provided in Appendix -A.
+In Section IV-A, we assess which of these probability distributions is the most appropriate for sparse training.
+III-B Dynamic sparse training
+Our complete dynamic sparse training algorithm starts by initializing each layer as a random bipartite graph using the Erdős–Rényi random graph generation algorithm [64].
+The number of active connections of the -th layer is given by , where is a parameter that controls the sparsity of the model [33].
 This initialization ensures that the number of active connections in a layer is proportional to its width, that is, .
-Moreover, Evci et al.
-(2020) showed that the Erdős–Rényi initialization also achieves better accuracy than uniform sparsity assignment.
-We verify their findings in Appendix F.
-The weights are initialized using the procedure proposed by Evci et al.
-(2019) which adapts standard weight initialization to take into account the actual connectivity of each unit and scales the weight distribution for each unit accordingly.
-Once the random sparse model is initialized, the weights of the active connections are optimized using stochastic gradient descent (SGD).
-Every training steps new connections are grown and the same number of connections are pruned.
+Moreover, the Erdős–Rényi initialization also achieves better accuracy than uniform sparsity assignment [35], as verified in Appendix -F.
+Once the initial connections are sampled, the weights are initialized so that the activations are normally distributed [65].
+During training, the weights of the active connections are optimized using stochastic gradient descent (SGD).
+Every training steps new connections are grown, which are added with a weight of zero, and the same number of connections are pruned.
 We provide pseudocode for our pruning and growing procedure in Algorithm 1, which operates on a single layer for simplicity.
@@
 This ensures that there is more exploration in the early stages of training.
-To determine the set of connections to grow , we start by evaluating one of the probability distributions described in Section 3.1.
+To determine the set of connections to grow , we start by evaluating one of the probability distributions described in Section III-A.
 Then, connections are sampled from the input and output distributions, where is a hyperparameter that adjusts the number of samples.
@@
 The grow set contains the largest gradient magnitude connections from the subset, where is the minimum between and to ensure that the grow set is not larger than the subset.
-The largest elements can be found in time using the introspective selection algorithm (Musser, 1997).
+The largest elements can be found in time using the introspective selection algorithm [66].
 The pruned set , in turn, contains the active connections with the smallest weight magnitude, also determined in time.
-Although magnitude pruning has been observed to cause layer collapse among pruning before training methods, Tanaka et al.
-(2020) investigated layer collapse and found that the training process is implicitly regularizing the weight magnitudes which minimizes the risk of layer collapse for DST methods like ours.
+Although magnitude pruning has been observed to cause layer collapse among pruning before training methods, the training process implicitly regularizes the weight magnitudes which minimizes the risk of layer collapse for DST methods like ours [32].
 Lastly, the active set and the weights are updated to reflect the changed connections.
-Pseudocode for the entire training process is provided in Appendix G.
+Pseudocode for the entire training process is provided in Appendix -G.
 It is important to stress that the dense model is not materialized at any point, only the active connections are represented, and all the operations are always sparse.
-4 Experiments
-At this point, in order to verify our claims, we need to determine what the minimum number of subset samples is that maintains the maximum accuracy.
-We answer this question in Section 4.2 and then compare the accuracy of our method against other sparsification methods in Section 4.3.
-We present accuracy results for ImageNet in Section 4.4.
-In Section 4.5, we explore the relation between model scale and accuracy for a fixed number of active connections by increasing the sparsity of larger models.
-Lastly, in Section 4.6, we compare the FLOPs of GSE with those of RigL.
-4.1 Experiment setup
-We evaluate our method and compare it to baselines on the CIFAR-10 and CIFAR-100 datasets (Krizhevsky et al., 2009), in addition to the ILSVRC-2012 ImageNet dataset (Deng et al., 2009).
-To establish how each method compares across model architectures we experiment with two convolutional neural networks (CNNs), ResNet-56 (He et al., 2016) and a 4 times downscaled version of VGG-16 (Simonyan and Zisserman, 2014), in addition to ViT (Dosovitskiy et al., 2020) modified for small datasets (Beyer et al., 2022) such that all models have roughly the same number of parameters.
+IV Experiments
+We evaluated our method and compared it to baselines on the CIFAR-10 and CIFAR-100 datasets [67], in addition to the ILSVRC-2012 ImageNet dataset [68].
+To establish how each method compares with different model architectures, we experimented with two convolutional neural networks (CNNs), ResNet-56 [69] and a 4 times downscaled version of VGG-16 [70], in addition to ViT [71] modified for small datasets [72] so that all models have roughly the same number of parameters.
 The bias and normalization parameters are kept dense since they only contribute marginally to the size and cost of a model.
-Similar to Wang et al.
-(2019), we evaluate all methods at 90%, 95%, and 98% sparsity.
+Like Wang et al.
+[41], we evaluated all methods at 90%, 95%, and 98% sparsity.
 The high levels of sparsity emphasize the differences between the methods.
-In addition, in Appendix L, we show that sparse matrix formats for unstructured sparsity become more efficient for matrix multiplication from 90% sparsity compared to dense matrix formats, further motivating the need for such high levels of sparsity.
-We repeat each experiment three times with random initializations and report the mean and plot the 95th percentile.
+In addition, in Appendix -K, we show that sparse matrix formats for unstructured sparsity become more efficient for matrix multiplication from 90% sparsity compared to dense matrix formats, further motivating the need for such high levels of sparsity.
+We repeat each experiment three times with random initializations.
+The plots report the mean and the 95th percentile, while the tables report the mean and standard deviation.
 We adopt the implementation of the baselines based on their available code.
-All experiments use the same optimization settings in order to isolate the differences in sparsification.
-Similar to Evci et al.
-(2020) and Lubana and Dick (2020), we use SGD with a momentum coefficient of 0.9, an regularization coefficient of 0.0001, and an initial learning rate of 0.1 which is dropped by a factor of 10 at epochs 80 and 120.
-The ViT experiments have one exception, they use an initial learning rate of 0.01.
+All experiments use the same optimization settings to isolate the differences in sparsification.
+Following Evci et al.
+[35] and Lubana et al.
+[57], we use SGD with a momentum coefficient of 0.9, an regularization coefficient of 0.0001, and an initial learning rate of 0.1 which is dropped by a factor of 10 at epochs 80 and 120.
+The ViT experiments have one exception: they use an initial learning rate of 0.01.
 We use a batch size of 128 and train for 200 epochs.
-We also apply standard data augmentation to the training data, described in Appendix D.
-Running all the reported experiments took on the order of 2000 GPU-hours with a single NVIDIA Tesla V100 16 GB GPU per training run on an internal cluster.
-The LTH procedure proposed by Frankle and Carbin (2018), denoted as Lottery in the results, uses iterative pruning with iterations at 0%, 50%, 75%, 90%, 95% and 98% sparsity.
-The gradual pruning method by Zhu and Gupta (2017), denoted as Gradual in the results, reaches the target sparsity at the second learning rate drop and prunes connections every 1000 steps.
+We also apply standard data augmentation to the training data, described in Appendix -D.
+The LTH procedure proposed by Frankle et al.
+[29], denoted as Lottery in the results, uses iterative pruning with iterations at 0%, 50%, 75%, 90%, 95% and 98% sparsity.
+The gradual pruning method by Zhu et al.
+[22], denoted as Gradual in the results, reaches the target sparsity at the second learning rate drop and prunes connections every 1000 steps.
 Following Tanaka et al.
-(2020), the pruning before training methods use a number of training samples equal to ten times the number of classes to prune the dense model to the target sparsity.
+[32], the pruning before training methods use a number of training samples equal to ten times the number of classes to prune the dense model to the target sparsity.
 Except for SynFlow, which uses 100 iterations with an exponential decaying sparsity schedule to prune the dense model.
-All the DST methods use the same update schedule: the connections are updated every 1000 steps, and similar to Evci et al.
-(2020), the fraction of pruned active connections is cosine annealed from 0.2 to 0.0 at the second learning rate drop.
+All DST methods use the same update schedule: the connections are updated every 1000 steps, and similar to Evci et al.
+[35], the fraction of pruned active connections is cosine annealed from 0.2 to 0.0 at the second learning rate drop.
 This is because Liu et al.
-(2021a) showed that DST methods struggle to recover from pruning when the learning rate is low.
+[50] showed that DST methods struggle to recover from pruning when the learning rate is low.
 The pruning and growing procedures are applied globally for all methods, maintaining the overall sparsity.
-This approach generally outperforms local, layer-wise sparsity (Mostafa and Wang, 2019).
-4.2 Effect of subset sample size
+This approach generally outperforms local, layer-wise sparsity [34].
+IV-A Effect of subset sample size
 First, we want to determine how the size of the subset affects the accuracy.
 In addition, we are interested in comparing the accuracy obtained by the three distributions: uniform, gradient upper bound (GraBo), and gradient estimate (GraEst).
-The number of subset samples (bounding the subset size) is set proportional to the number of active connections , with and .
+The number of subset samples (which is a bound on the subset size) is set proportional to the number of active connections , with and .
 The results are shown in the first three columns of Figure 3 and include the baseline accuracy of RigL for comparison.
@@
 The observations align precisely with our expectations, which further underscores that greedy exploration leads to inferior solutions when compared to the hybrid exploration strategy of GSE.
-Interestingly, uniform sampling demonstrates strong performance while we hypothesized that the biased distributions would be beneficial for training.
-This shows that exploring connections which, at first, do not seem beneficial for training, can in fact be quite helpful.
-It would be interesting for future research to investigate exactly which training dynamics are enabled by uniform sampling.
-Since uniform sampling is also the most efficient distribution, we use it as our method in the following experiments.
-4.3 Comparison with related work
-We compare our method to a broad range of sparsification methods in Table 1, the best accuracy among the sparse training methods is bolded.
-The Static method denotes training a static random sparse model.
-The other baselines are Lottery (Frankle and Carbin, 2018), Gradual (Zhu and Gupta, 2017), SNIP (Lee et al., 2018), GraSP (Wang et al., 2019), SynFlow (Tanaka et al., 2020), SET (Mocanu et al., 2018), RigL (Evci et al., 2020), and Top-KAST (Jayakumar et al., 2020).
-Our results use GSE with uniform sampling and .
 | Dataset | CIFAR-10 | CIFAR-100 | ||||||||||||||||
@@
 | GSE (ours) | 91.0 | 89.9 | 87.0 | 80.0 | 76.4 | 70.7 | 88.6 | 88.1 | 85.1 | 66.0 | 62.6 | 54.3 | 54.4 | 48.8 | 40.7 | 57.9 | 52.3 | 40.8 |
-Among the sparse training methods (from Static downwards in Table 1) our method outperforms the other methods consistently over datasets, sparsities, and model architectures.
-While at 90% sparsity all the methods achieve comparable accuracy, at the extreme sparsity rate of 98%, differences between methods become more evident.
+Surprisingly, uniform sampling demonstrates strong performance while we hypothesized that biased distributions would be beneficial for training.
+This shows that exploring connections which, at first, do not seem beneficial for training, can in fact be quite helpful.
+It would be interesting for future research to investigate exactly which training dynamics are enabled by uniform sampling.
+Since uniform sampling is also the most efficient distribution, we use it as our method in the following experiments.
+IV-B Comparison with related work
+We compare our method with a wide range of sparsification methods in Table I on CIFAR-10 and CIFAR-100.
+The best accuracy among the sparse training methods is in bold.
+The Static method denotes training a static random sparse model.
+The other baselines are Lottery [29], Gradual [22], SNIP [31], GraSP [41], SynFlow [32], SET [33], RigL [35], and Top-KAST [36].
+Our results use GSE with uniform sampling and .
+Among the sparse training methods (from Static downwards in Table I) our method outperforms the other methods consistently over datasets, sparsities, and model architectures.
+While at 90% sparsity all methods achieve comparable accuracy, at the extreme sparsity rate of 98%, differences between methods become more evident.
 At 98% sparsity our method outperforms all the before training sparsification methods by a significant margin, 4.7% on average.
-GSE also consistently matches or surpasses both RigL and Top-KAST, while being more efficient.
+GSE also consistently matches or exceeds both RigL and Top-KAST, while being strictly more efficient.
 For example, with VGG-16 on CIFAR-10, GSE achieves 1.4% and 4.6% higher accuracy than RigL and Top-KAST, respectively.
@@
 However, this comes at the cost of training the dense model, thus requiring significantly more FLOPs than our method.
-This observation serves as motivation for further research into dynamic sparse training methods that improve on our accuracy while preserving our efficiency advantage.
+This observation serves as motivation for further research into dynamic sparse training methods that improve our accuracy while preserving our efficiency advantage.
 Notably, our method outperforms Lottery on the CNN models, while Lottery requires multiple rounds of training.
-In Appendix H, we report the test accuracy of each method at every epoch throughout training, showing that our method trains as fast as the other sparse training methods.
-Additionally, in Appendix M, we provide plots of the layer-wise sparsity at the end of training for each combination of method, model, and sparsity.
-This information provides insight into how each method distributes the available connections over the layers of a model.
-4.4 ImageNet
+In Appendix -H, we report the test accuracy of each method at every epoch throughout training, showing that our method trains as fast as the other sparse training methods.
+IV-C ImageNet
+To determine the efficacy and robustness of GSE on large models and datasets, in this third experiment, we compare the accuracy of ResNet-50 trained on ImageNet at 90% sparsity.
+In Table II, the accuracy obtained with GSE is compared against the baselines from the preceding section, with the additions of DSR [34] and SNFS [61].
+Since this is a common benchmark among sparse training methods, the accuracy results of the baselines were obtained from prior publications.
 | Method | Accuracy |
@@
 | GSE (ours) | 73.2±0.07 |
-To determine the efficacy and robustness of GSE on large models and datasets, in this third experiment, we compare the accuracy of ResNet-50 trained on ImageNet at 90% sparsity.
-In Table 2, the accuracy obtained with GSE is compared against the baselines from the preceding section, with the additions of DSR (Mostafa and Wang, 2019) and SNFS (Dettmers and Zettlemoyer, 2019).
-Since this is a common benchmark among sparse training methods, the accuracy results of the baselines were obtained from prior publications.
 Our experiments use the aforementioned hyperparameters with the following exceptions: we train for 100 epochs and use learning rate warm-up over the initial 5 epochs, gradually reaching a value of 0.1; subsequently, the learning rate is reduced by a factor of 10 every 30 epochs; and we use label smoothing with a coefficient of 0.1.
 Our results use GSE with uniform sampling and .
-The results in Table 2 align with those for CIFAR, and further underscore GSE’s superior performance over other sparse training methods.
-Note also that GSE again outperforms less efficient methods such as SNFS, RigL, and Top-KAST.
-These results collectively reinforce that GSE is robust and able to scale to large models and datasets.
-4.5 Model scaling
+The results in Table II show a 0.2% improvement in model accuracy, which is the largest increase since the introduction of SNFS [61] in 2019 and sets a new state-of-the-art in dynamic sparse training.
+These results align with those for CIFAR, and further underscore GSE’s performance over other sparse training methods.
+Note that GSE not only obtains a higher accuracy, it is also more efficient than methods such as SNFS, RigL, and Top-KAST.
+These results collectively reinforce GSE’s robustness and its ability to scale to large models and datasets.
+IV-D Model scaling
 In this experiment, we investigate how the size of a model affects the accuracy when the number of active connections is kept constant.
-We scale the number of filters in convolution layers and units in fully-connected layers to obtain a wider model.
-The results, presented in Figure 3, show that training wider CNN models translates to significantly increased accuracy, 6.5% on average at 4 times wider.
-These findings support those from Zhu and Gupta (2017) and Kalchbrenner et al.
-(2018), however, our method enables a sparse model to be trained whose dense version would be too computationally intensive.
+We scaled the number of filters in convolution layers and units in feedforward layers to obtain a wider model.
+The results, presented in Table III, show that training wider CNN models translates to significantly higher accuracy, 6.5% on average at 4 times wider.
+These findings support those of Zhu et al.
+[22] and Kalchbrenner et al.
+[23], however, our method enables a sparse model to be trained whose dense version would be too computationally intensive.
+| Width / Sparsity | 1x / 0% | 2x / 75% | 3x / 89% | 4x / 94% |
+| ResNet-56 | 70.1±0.4 | 72.3±0.4 | 73.8±0.6 | 74.2±0.3 |
+| VGG-16 | 62.3±0.8 | 68.4±0.7 | 70.4±0.8 | 71.2±0.3 |
+| ViT | 58.8±0.5 | 58.3±0.2 | 57.3±0.7 | 56.6±0.1 |
 The same trend does not hold for ViT, as its accuracy decreases with increasing scale.
-Appendix J reveals that the accuracy increases slower for larger ViT models, while for CNNs the accuracy increases faster as they get larger.
-Moreover, the accuracy of larger dense ViT models plateaus at 62.2%, as shown in Appendix K, indicating that this phenomenon is independent of our method.
-ViT models are known to require large amounts of training data to achieve optimal accuracy (Dosovitskiy et al., 2020), we conjecture that this contributes to the trend observed with ViT.
-In Appendix I, we observe that both ViT and CNN models do benefit considerably from extended training, with a 5.3% average increase in accuracy for training twice as long.
-4.6 Floating-point operations
-Lastly, to determine the efficiency of GSE, we compare its FLOPs to those of RigL in Figure 4.
-The FLOPs are obtained for ResNet-50 on ImageNet using the method described by Evci et al.
-(2020), with .
-The results show that at 99% sparsity GSE uses 11.8% fewer FLOPs than RigL, at higher sparsities GSE uses considerably fewer FLOPs.
-For illustration, the right plot extends the sparsity to that of various animal brains which underscores the significance of our always sparse training method, as RigL uses times more FLOPs to train a model with the same sparsity as the cat brain.
-4.7 Limitations
-While we experiment on a range of sparsities, datasets, models, and baseline methods, we acknowledge that further exploration would enhance the robustness of our findings.
-Following previous work, our experiments focus on the domain of computer vision, and although we demonstrate the effectiveness of our approach within this domain, it remains to be seen how our results generalize to other domains such as natural language processing and reinforcement learning, among others.
-Despite this, we believe that our proposed always-sparse training algorithm, supported by its linear time complexity with respect to the model width during training and inference, represents a significant advancement.
-Furthermore, our method outperforms existing sparse training algorithms in terms of accuracy, as demonstrated on the CIFAR and ImageNet datasets using ResNet, VGG, and ViT models, alongside a comprehensive comparison with a range of sparsification techniques.
-Secondly, our accuracy results were obtained by simulating sparse networks using binary masks.
-This is standard practice among sparse training methods because support for sparsity in machine learning frameworks is still in the early days, making it challenging to train models using sparse matrix formats for the weights.
-We note that there is nothing about our method that prevents it from being accelerated on GPUs, yet the significant engineering required to develop an optimized GPU kernel falls outside the scope of this work.
-To ensure that our claims are still validated, we compare the FLOPs of our method with those of RigL in Section 4.6.
-Moreover, we present an unstructured sparse matrix multiplication benchmark in Appendix L, which shows that a 10 times reduction in execution time is achieved at 99% sparsity, compared to dense matrix multiplication.
-Thus, evidently showing that real speedups can be achieved, even with unstructured sparsity.
-5 Conclusion
+Appendix -I reveals that the accuracy increases slower for larger ViT models, while for CNNs the accuracy increases faster as they get larger.
+Moreover, the accuracy of larger dense ViT models plateaus at 62.2%, as shown in Appendix -J, indicating that this phenomenon is independent of our method.
+ViT models are known to require large amounts of training data to achieve optimal accuracy [71], we conjecture that this contributes to the trend observed with ViT.
+IV-E Extended training
+We continued our experimentation by looking at the effect of extended training.
+The following extended training results are for CIFAR-100 at 98% sparsity with 1x, 1.5x, and 2x the number of training epochs.
+The learning rate drops and the final prune-grow step are also scaled by the same amount, all other hyperparameters stay the same.
+We observe that ViT and CNN models benefit considerably from extended training, with an average increase in accuracy of 5.3% for training twice as long.
+The ViT model in particular gains 7% in accuracy by training twice as long.
+| Epochs | 200 | 300 | 400 |
+| ResNet-56 | 54.3±0.5 | 56.4±0.3 | 57.9±0.1 |
+| VGG-16 | 40.8±0.3 | 44.1±0.4 | 45.9±0.2 |
+| ViT | 40.7±1.0 | 45.3±1.1 | 47.8±0.5 |
+IV-F Floating-point operations
+Lastly, to determine the training efficiency of GSE, we compare its FLOPs with those of RigL in Figure 4.
+FLOPs are determined for ResNet-50 on ImageNet using the method described by Evci et al.
+[35], with .
+The results show that the benefit of GSE increases rapidly at high sparsity levels.
+At 99% sparsity, GSE uses 13.3% fewer FLOPs than RigL, up from 2.4% at 95% sparsity.
+V Conclusion
 We presented an always-sparse training algorithm using guided stochastic exploration of the connections which has a linear time complexity with respect to the model width and improves upon the accuracy of previous sparse training methods.
@@
 Interestingly, we showed that training larger and sparser CNN models results in higher accuracy for the same number of active connections.
+Moreover, it was observed that our method greatly benefits from extended training.
 Lastly, the number of floating-point operations of our method decrease considerably faster than those for RigL as the sparsity increases.
 This observation underscores the efficiency and scalability of our method, which is promising for training increasingly larger and sparser ANNs.
-Appendix A Connection sampling distributions
-In this section we provide a more detailed discussion on the distributions which were considered for sampling the subset of inactive connections.
-A.1 Gradient magnitude upper bound (GraBo)
-The gradient magnitude for a single connection can be expressed as follows:
-where and are the activations and gradients of the loss at the output units of the -th layer, respectively.
-When the batch size is one, the sum disappears, simplifying the gradient magnitude to a rank one matrix which can be used to sample connections efficiently, inducing a joint probability distribution that is proportional to the gradient magnitude only when the batch size is one.
-In practice, however, training samples come in mini batches to reduce the variance of the gradient.
-Therefore, we experiment with sampling the connections proportionally to the following upper bound of the gradient magnitude instead, which enables efficient sampling, even with mini batches:
-The proof for Equation 2 uses the triangle inequality and is provided in Appendix B.
-This upper bound has the property that it becomes an equality when .
-Connections are then sampled from the probability distributions in Equation 3.
-The implementation of this distribution does not require any modifications to the model, and only needs a single forward and backward pass to evaluate.
-A.2 Gradient magnitude estimate (GraEst)
-The reason we use a bound on the absolute gradient is that in general the sum of products cannot be written as the product of sums, that is, .
-However, it is possible to satisfy this equality in expectation.
-This was shown by Alon et al.
-(1996) who used it to estimate the second frequency moment, or the norm of a vector.
-Their work is called the AMS sketch, AMS is an initialism derived from the last names of the authors.
-They then showed that this also works in the more general case to estimate inner products between vectors (Alon et al., 1999).
-This is achieved by assigning each index of the vector a random sign, i.e.
-, resulting in the following equality:
-We can use their finding to construct probability distributions and in Equation 5 such that their induced joint probability distribution is proportional to the gradient magnitude in expectation.
-This distribution can also be computed in a single forward and backward pass and requires no modifications to the model, apart from introducing a mechanism to sample the random signs .
-Appendix B Gradient magnitude upper bound proof
-Lemma B.1.
-The gradient magnitude of the loss with respect to the parameters of a fully-connected layer has the following upper bound:
-where , , and are the weight matrix, activations and gradients of the loss at the output units of the -th layer, respectively.
-The number of units in the -th layer is denoted by , and is the batch size.
-Proof.
-The proof starts with the definition of the gradient of the fully-connected layer, which is calculated during back propagation as the activations of the previous layer multiplied with the back propagated gradient of the output units of the layer.
-We then write this definition for a single connection in the weight matrix and use the triangle inequality to specify the first upper bound.
-In Equation 8 we rewrite the upper bound of Equation 7 as all the cross products between the batch dimensions minus all the non-identical cross terms.
-Note that all the non-identical cross terms are positive because of the absolute, therefore we get another upper bound by removing the non-identical cross terms which we then write in matrix notation in Equation 9.
-Appendix C Data normalization
-We normalize all the training and test data to have zero mean and a standard deviation of one.
-The dataset statistics that we used are specified below, where is the mean and the standard deviation.
-The values correspond to the red, green, and blue color channels of the image, respectively.
-CIFAR-10
-CIFAR-100
-ImageNet
-Appendix D Data augmentation
-We applied standard data augmentation as part of our training data pipeline.
-The specific augmentation techniques are specified below.
-We apply the data normalization described in Appendix C after the augmentation transformations.
-CIFAR-10
-We pad the image with 4 black pixels on all sides and randomly crop a square of 32 by 32 pixels. We then perform a random horizontal flip of the crop with a probability of 0.5.
-CIFAR-100
-We pad the image with 4 black pixels on all sides and randomly crop a square of 32 by 32 pixels. We then perform a random horizontal flip of the crop with a probability of 0.5.
-ImageNet
-We randomly crop a square of 224 by 224 pixels.
-We then perform a random horizontal flip of the crop with a probability of 0.5.
-For the test set, we resize the images to be 256 pixels followed by a square center crop of 224 pixels.
-Appendix E Convolutional layer subset sampling
-In this section we describe how our growing and pruning algorithm can be applied to convolutional layers.
-While a weight in a fully-connected layer is specified by the input and output units that it connects, a weight in a 2D convolutional layer is specified by the input channel, its and coordinate on the filter, and the output channel.
-It thus seems that it would require four discrete probability distributions to sample a weight of a 2D convolutional layer, instead of two for fully-connected layers.
-However, one can alternatively view a convolutional layer as applying a fully-connected layer on multiple subsections (patches) of the input, flattening the input channels and filters makes the weights of a convolutional layer identical to a fully-connected layer.
-The multiple subsections of the input can then simply be treated as additional batches.
-With this realization all the methods that we explain throughout the paper for fully-connected layers directly apply to convolutional layers.
-Appendix F Sparse graph initialization comparison
-In Figure 5, we plot the accuracy obtained by assigning the sparsity uniformly over all the layers compared to using the Erdős–Rényi initialization for SET, RigL, and our method.
-Note that Ours Uniform (Erdős–Rényi) denotes uniform sampling of the subset connections and the Erdős–Rényi initialization.
-The Erdős–Rényi initialization obtains better accuracy while ensuring that the number of active connections per layer is proportional to the layer width.
-Appendix G Pseudocode
-We provide pseudocode for our complete dynamic sparse training process in the case of the GraBo or GraEst distributions in Algorithm 2.
-The uniform distribution is simpler as it does not require any aggregation of activity.
-For simplification, the pseudocode assumes that the prune-grow step only uses a single batch of training samples, allowing the aggregation for the probability distributions f and g to be computed in the same pass as the generation of the various connection sets.
-This is also the scenario we tested with in all our experiments.
-Appendix H Training progress
-In Figure 6, we provide plots for the test accuracy during training for each baseline method and our method.
-It can be seen that our method consistently achieves among the highest test accuracy throughout training compared to the sparse training methods.
-The training progress of Gradual stands out because its accuracy increases the most at first, this is because it starts out as a dense model while it gradually increases the model sparsity, as a result, the accuracy goes down towards the second learning rate drop when the target sparsity is reached.
-Appendix I Extended training
-The following extended training results are for CIFAR-100 at 98% sparsity with 1x, 1.5x, and 2x the number of training epochs.
-The learning rate drops and the final prune-grow step are also scaled by the same amount, all other hyperparameters stay the same.
-It is clear that all models benefit significantly from extended training.
-The ViT model in particular gains 7% in accuracy by training for twice as long.
-| Epochs | 200 (1x) | 300 (1.5x) | 400 (2x) |
-| ResNet-56 | 54.3±0.5 | 56.4±0.3 | 57.9±0.1 |
-| VGG-16 | 40.8±0.3 | 44.1±0.4 | 45.9±0.2 |
-| ViT | 40.7±1.0 | 45.3±1.1 | 47.8±0.5 |
-Appendix J Model scaling training progress
-In Figure 7, we provide plots for the test and training accuracy during training for our method on CIFAR-100 while increasing the width of the model by 1x, 2x, 3x, and 4x.
-The number of active connections is kept constant between all experiments, see Section 4.5 for details.
-In the CNN models, ResNet-56 and VGG-16, the increase in width causes the models to train significantly faster and obtain higher accuracy.
-For the vision transformer (ViT) model, the trend is reversed, wider models train slower and obtain lower accuracy.
-Appendix K Scaling dense ViT
-To further investigate the behavior of ViT, which decreases in accuracy when using larger models with the same number of activate parameters, as discussed in Section 4.5 and Appendix J.
-In Table 4, we report the accuracy of ViT for increasing model width but kept dense, thus increasing the active connections count.
-We see the accuracy quickly plateaus at 62.2%, gaining only 2% accuracy while even the sparse CNNs gained 6.5% on average at 4 times wider.
-This indicates that the decreasing accuracy of ViT reported in Section 4.5 is not due to our sparse training method as even the dense model does not benefit much from increased scale.
-| Model width | 1x | 2x | 3x | 4x |
-| Accuracy | 60.3±0.6 | 62.5±0.5 | 62.2±0.4 | 62.2±0.1 |
-Appendix L Sparse matrix multiply benchmark
-In Figure 8, we show the execution time of the sparse times dense matrix-matrix multiply on two execution devices: Intel Xeon Gold 6148 CPU at 2.40GHz and NVIDIA Tesla V100 16GB.
-The sparse matrix is a square matrix of size units times units, representing an unstructured sparse random weight matrix.
-The dense matrix is of size units times 128 and represents a batch of layer inputs.
-We compare three storage formats, namely: dense, where all the zeros are explicitly represented; Coordinate format (COO), which represents only the nonzero elements and their coordinates; and Compressed Sparse Row (CSR), which is similar to COO but compresses the row indices.
-Each data point in Figure 8 is the average of 10 measurements.
-The results show that the CSR format generally outperforms the COO format.
-Moreover, the execution time of the CSR format improves upon the dense execution time starting around 90% sparsity.
-At 99% sparsity, the CSR format is up to an order of magnitude faster than the dense format.
-These results are similar between the two execution devices, although the NVIDIA Tesla V100 16GB is up to two orders of magnitude faster than the Intel Xeon Gold 6148 CPU at 2.40GHz.
-Appendix M Layer-wise sparsity
-In Figure 9, we show the sparsity of each layer, in terms of compression, obtained by each sparsification methods at the end of training.
-In the plots the CIFAR-10 and CIFAR-100 results are averaged.
-These plots can be used to gain insight into how each sparsification method distributes the active connections over the models.